A Bibliography of Publications about the AT&T Plan 9 Distributed Operating System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 October 2024 Version 1.50

Title word cross-reference

81/2 [Pik91, Pik00a]. \frac{81}{2} [Pikxxa]. α [PT00]. ϵ [PT00, PT00]. η [PT00]. κ [PT00]. λ [PT00]. μ [PT00]. ρ [PT00]. σ [PT00].

11th [USE02]. 18th [IEE04]. 19th [ABB93].

5 [Ano96a].

64-bit [TCxx].

9 [Ano90, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano96d, Ano96c, Anoxxa, Anoxxb, Anoxxc, Anoxxd, Ano00a, Bai96b, Bai96a, BGSL05, Cox00, CGP^{+}02, DPP92, Duf90a, Duf90b, Dufxx, Duf00, Fen96, Fil96, Fla00, Fla02, Gan08, Goo90, GS93, HF00, Kor96, Kot94, Kue97, Lug81, MKM^{+}14, MS04, PPTT90a, PPTT90b, PPTT91a, PPT91a, PPT^{+}91b, Pik91, PPT^{+}92, PPT^{+}93, PPD^{+}95, P^{+}96a, P^{+}96b, Pikxxb, PPT^{+}xx, Pikxxa, Pik00c, Pik00d, Pik00a, Pik02, Pre88, Pre90, PPTT91b, PPTT92, Pre93a, Pre93b, PW93, PPTT94, Prexx, PPTTxxb, Ray94, Sko98, Sta12, TCxx, Thoxxb, Tho00b, Wel93, Wel94, WJV^{+}84]. ’93 [ABB93]. 9p [MM06]. 9p-based [MM06].

Acid [Win00a, Win94, Win00b]. Acme [Pik94]. Adding [Fla02]. aggregate [VFMM08]. algorithm [Tho68]. Algorithms [KP99b]. Analytical [JCSM96, JCMS97]. Annual [USE00]. ANSI [Tri00]. ANSI/POSIX [Tri00]. APE [Tri00]. Application [Fla02, Lau00]. Applications [BGSL05]. Approach [Cox09, MS04, QD02]. April [IEE04, USE92]. Architecture [Fla02]. Architectures [USE92, Wel93, Wel94]. archival [QD02]. assembler [Pik00e]. AT&T [Ano95a, Ano96b, Bai96b, Bai96a]. August [ABB93, USE02]. authentication [Gan08]. Automata [MY60, Lau00, RS59].

Bad [Fen96]. Bangs [MKM^{+}14]. based [Ier09, MM06, Win94]. bases [ABB93]. Be [Cox07]. Beautiful [OW07]. Becomes [BGSL05]. Bell [PPTTxxa, BGSL05, PPTT90a, PPTT90b, PPD^{+}95, Pre88, Ray94, Trb16]. Beyond [Ano90, Ano96a]. Big [MKM^{+}14]. bit [TCxx]. Brace [Ano96a]. Built [Win00a].

C [Pikxxb, Pik00b, Thoxxa, Thoxxb, Tho00a]. CA [USE00, USE02]. Cached [Qui91]. California [USE93, USE94]. Came [Trb16]. Can [BGSL05, Cox07]. Case [BGSL05]. Changes [Pik02]. clusters [MM06]. code [OW07, Cox12]. comment [JCMS97]. Communications [IEE05]. comparison [Wel93, Wel94]. Compiler [Pikxxb, Pik00b, Thoxxa]. Compilers [Thoxxb, Tho00a]. Compressed [ZMSD93]. computer [PPT91a]. Computing [Ano96b, IEE05, MKM^{+}14, MS04]. Conference [ABB93, Eur91, IEE05, USE88, USE93, USE94, USE00]. contrary [PPT91a]. controls [Ano95a]. conversion [Lau00]. Costs [Ano96a]. Cutting [Ano96a].

data [ABB93]. Databases [Ano96a]. Dead [MI07]. Debugger [Win94, Win00a]. decision [RS59]. Designing [JCSM96, PPTT91a, PPT91a, JCMS97]. Desktop [Ano96a]. deterministic [Lau00]. Diego [USE93, USE00]. Distributed [Ano96b, BJ94, Eur91, IEE04, MKM^{+}14, PPTT94, PPTTxxb, PPT91a, PPTT91b, PPTT92, Ray94, Wel93, Wel94].

Distribution

[Anoxxa, Anoxxc, Anoxxd, Ano00a]. DMI [Ano96a]. Documents [P^{+}96b]. Dot [Pik00c, Pik00d]. Dot-Dot [Pik00c, Pik00d]. Downsizing [Ano96a]. Draft [Kot94]. Dublin [ABB93].

Earlier [Sta12]. Early

[PPT^{+}91b, PPT^{+}xx]. Edge [Ano96a]. Editor [Pik87, Pik00h, Ritxx, DL67]. efficient [Cox10b]. England [USE88]. Enumerating [McI04]. Environment [Pik02, Tri00, VFMM08]. Errata [Anoxxd]. EurOpen [Eur91]. EUUG [USE88]. Everywhere [Ano96a]. expect [PPT91a]. Expression [Cox07, Cox09, Cox10a, Cox12, Ier09, Cox10b, Ker07, Tho68]. Expressions [KP99b, MY60, Lau00].

Fast [Cox07]. Fe [IEE04]. Feature [Fil96]. Feature-Rich [Fil96]. Fi [Goo90]. File [Pik00c, Pik00d, Qui91, TCxx, Tho00b, Wel93, Wel94]. Files [HF00, ZMSD93]. Film [Fil96]. Finite [RS59]. First [Ano96a]. Fourth [Pik02]. Francisco [USE94, USE02]. Future [Goo90].

Generation [Kue97]. gets [GS93]. Getting [Cox00, Pik00c, Pik00d]. GNU [Ano95b, Ano95c, Rob95a, Rob95b]. Go [Ano96a]. Good [Fen96]. Google [Cox12]. Grammars [Ier09]. Graphs [MY60]. Grid [MS04]. grids [MM06]. Guide [Gli00, Kot94, GS93].

Hawaii [IEE05]. Hello [PT93, PT00]. Hierarchies [JCSM96, JCMS97]. Hike [Ano96a]. history [Ritxx]. hold [Ano95a]. Holistic [VFMM08].

IEEE [IEE05]. II [Rob95b]. incomplete [Ritxx]. Index [Cox12]. Inferno [Ano96a]. installation [GS93]. Installing [Anoxxa, Ano00a]. integrated [MS04].

Interface [Pik94]. International [ABB93, IEE04, IEE05]. Inverted [ZMSD93]. Ireland [ABB93]. Irrelevant [Pik00g]. ISDN [Ano96a]. Island [IEE05].

January [USE93, USE94]. Java [Ano96a]. June [USE00].

Kauai [IEE05]. Kernel [Ano96c, Kor96, USE92]. Kernels [USE92]. Konzepte [B^{+}99].

Laboratories [Ray94]. Labs [BGSL05, PPTT90a, PPTT90b, PPD^{+}95, PPTTxxa, Pre88, Trb16]. Language [Luc97, Win94, Win00a]. Languages [KP99b, McI04]. Large [ZMSD93, ABB93]. Lexical [Pik00c, Pik00d]. Lexicons [ZMSD93]. library [Cox10b]. License [Sta12]. like [PPT91a]. Limbo [Luc97]. line [Ano95a]. Link [Ano96a]. Linux [Gan08, Min01]. List [Anoxxd]. London [USE88]. look [PPT91a]. lot [PPT91a]. lp [Gli00].

Machine [Cox09]. maintain [Ano95a]. Maintaining [HF00]. man [GS93]. management [MM06]. Manual [OK00, P^{+}96a, P^{+}96b, Pik00e, Win00b]. Manuals [P^{+}96a]. March [IEE05]. matcher [Ker07]. Matching [Cox07, Cox09, Cox10a, Cox12, Ier09]. Matter [MI07]. May [Eur91, PPT91a]. media [Ano96d]. Memory [JCSM96, JCMS97, PPTHxx, PPTH00]. Mexico [IEE04]. Micro [USE92]. Micro-Kernels [USE92]. might [PPT91a]. Mk [HF00]. Mkfiles [Fla00]. Mobile [Ano96d]. Model [JCSM96, JCMS97]. modifications [Ano95a]. Movie [Fen96]. Multi [Ano96d]. Multi-media [Ano96d]. Multics [MKM^{+}14]. Multiprocessor [PPTHxx, PPTH00, Pre90, Pre93a, Pre93b, Prexx]. naechste [Kue97]. Name [PPT^{+}92, PPT^{+}93]. Names [Pik00c, Pik00d]. Namespaces [Min01]. Network [Pre88]. networks [PW93]. Netzbetriebssystem [B^{+}99]. News [Kor96]. NFAs [Lau00]. Norway [Eur91].

o [PT00]. Object [Ano96a]. online [DL67]. Open [Eur91, BJ94]. Operating [Fen96, Ano95a, Ray94]. organization [PW93]. OSes [MI07]. Other [Pik00f, USE92]. Outer [Ano96c, Kor96, Lug81, WJV^{+}84].

Papers [PPT^{+}91b, PPT^{+}xx]. Parallel [IEE04]. Parsing [Ier09]. Part [Rob95a, Rob95b, Ano95b, Ano95c]. Partially [ZMSD93]. pattern [Ier09]. pattern-matching [Ier09]. PC [Bai96a]. PerCom [IEE05]. Perspective [Eur91]. Pervasive [BGSL05, IEE05]. Plan [Ano96b, B^{+}99, Kue97, Pik00e, Sko98, Ano90, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano96d, Ano96c, Anoxxa, Anoxxb, Anoxxc, Anoxxd, Ano00a, Bai96b, Bai96a, BGSL05, Cox00, CGP^{+}02, DPP92, Duf90a, Duf90b, Dufxx, Duf00, Fen96, Fil96, Fla00, Fla02, Gan08, Goo90, GS93, HF00, Kor96, Kot94, Lug81, MKM^{+}14, MS04, PPTT90a, PPTT90b, PPTT91a, PPT91a, PPT^{+}91b, Pik91, PPT^{+}92, PPT^{+}93, PPD^{+}95, P^{+}96a, P^{+}96b, Pikxxb, PPTTxxa, PPT^{+}xx, Pikxxa, Pik00b, Pik00c, Pik00d, Pik00a, Pik02, Pre88, Pre90, PPTT91b, PPTT92, Pre93a, Pre93b, PW93, PPTT94, Prexx, PPTTxxb, Ray94, Rob95a, Rob95b, Sta12, TCxx, Thoxxb, Tho00a, Tho00b, Wel93, Wel94, WJV^{+}84]. plans [Ano95a]. Plumbing [Pik00f]. poor [GS93]. Ports [Ano00b]. POSIX [Sko98, Tri00]. Practice [KP99a]. Price [Ano96a]. principled [Cox10b]. Printer [Gli00]. Private [Min01]. Problems [Sta12, RS59]. Proceedings [Eur91, USE88, USE92, USE93, USE94, USE02, ABB93, IEE05, IEE04]. Process [PPTHxx, PPTH00, MM06]. Processing [IEE04]. Professionals [Ano96a]. Programmer [P^{+}96a, P^{+}96b]. Programmers [Pik94]. Programmierung [B^{+}99]. Programming [KP99a, Luc97, Pik02, Tho68]. protocol [PW00].

QED [Ritxx].

RAID [Ano96a]. RAID-5 [Ano96a]. Rc [Duf90a, Duf90b, Dufxx, Duf00]. re2 [Cox10b]. Reduce [Ano96a]. Regular [Cox07, Cox09, Cox10a, Cox12, KP99b, MY60, Cox10b, Ker07, Lau00, McI04, Tho68]. Relational [Ano96a]. Release [Pik02]. Research [Pik00g]. resource [VFMM08]. Rich [Fil96]. Right [Pik00c, Pik00d].

sam [Pik87, Pik00h]. San [USE93, USE94, USE00, USE02]. Santa [IEE04]. Scalable [Ano96b]. School [Ano96a]. Sci [Goo90]. Sci-Fi [Goo90]. Search [Cox12, Tho68]. Searching [ZMSD93]. Seattle [USE92]. Security [CGP^{+}02, USE02, Ray94]. Server [TCxx, Tho00b]. Service [Ano96a]. Setup [Kot94]. Shared [PPTHxx, PPTH00]. Shared-memory [PPTHxx, PPTH00]. Shell [Dufxx, Duf00, Duf90a, Duf90b]. Simple [Cox07]. Sleep [PPTHxx, PPTH00]. Slows [Ano96a]. Software [KP99b, Pik00g]. Some [MI07]. Space [Ano96c, Kor96, Lug81, WJV^{+}84]. Spaces [PPT^{+}92, PPT^{+}93]. Specified [ZMSD93]. SPIN [Hol00]. Spooler [Gli00]. Spring [Eur91, USE88]. Sprite [Wel93, Wel94]. Standalone [TCxx]. Started [Cox00]. State [MY60]. Still [MI07]. storage [QD02]. Streams [Pre90, Pre93b, Prexx, Pre93a]. strings [McI04]. Study [BGSL05]. Summer [Ano96a]. Support [Ano96a, Fla02].

Symposium [IEE04, USE02]. System [Fen96, Kot94, MKM^{+}14, Pikxxa, Pik00g, PPTT94, PPTTxxb, Qui91, Ray94, Ano95a, MM06, Pik91, Pik00a, PPTT91b, PPTT92, Wel93, Wel94]. Systems [BGSL05, Dufxx, Eur91, BJ94, Duf90a, Duf90b, PPT91a].

tagged [Lau00]. Technical [USE00]. Techniques [Tho68]. Terms [ZMSD93]. Text [Pik87, Pik00h, Ritxx, Ier09]. their [Lau00, RS59]. there [GS93]. Third [IEE05]. three [Wel93, Wel94]. tight [Ano95a]. tomorrow [PPT91a]. tool [Ier09]. Traditional [BGSL05]. transitions [Lau00]. Trigram [Cox12]. Troff [OK00]. Tromsø [Eur91].

Ubiquitous [BGSL05]. ugly [GS93].

Universe [MKM^{+}14]. UNIX

[Bai96a, Dufxx, Eur91, Goo90, Sko98, Ano90, Duf90a, Duf90b]. USA [USE92, USE93, USE94, USE00]. Use [Ano96a, PPT^{+}92, PPT^{+}93, Pikxxb, Pik00b]. User [OK00, Pik94]. Using [Hol00, ZMSD93]. Utilities [Pik00f].

Various [Ano00b]. Venti [QD02]. Very [ABB93]. Virtual [Cox09]. VLDB [ABB93]. Vnode [Wel93, Wel94]. Volume [P^{+}96a, P^{+}96b].

WA [USE92]. Wakeup [PPTHxx, PPTH00]. Well [BGSL05]. Wild [Cox10a]. Window [Pikxxa, Pik91, Pik00a]. Winter [USE93, USE94]. Work [BGSL05]. worked [Cox12]. Workshop [USE92]. Workshops [IEE05]. World [PT93, PT00]. WORM [Qui91].

XCPU [MM06].

Yourself [Ano96a].

References

Agrawal:1993:VLD

[ABB93] Rakesh Agrawal, Sean Baker, and David Bell, editors. Very large data bases, VLDB ’93: proceedings of the 19th International Conference on Very Large Data Bases, August 24–27, 1993, Dublin, Ireland. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1993. ISBN 1-55860-152-X. LCCN QA76.9.D3 I61 1993. Co-sponsored by VLDB Endowment and Irish Computer Society; in co-operation with the IEEE Technical Committee on Data Engineering.

Anonymous:1990:BUP

[Ano90] Anonymous. Beyond Unix: Plan 9. Information Week, 293:20–??, October 29, 1990. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1995:APH

[Ano95a] Anonymous. AT&T plans to hold the line and maintain tight controls on modifications to its new Plan 9 operating system. PC Week, 12(30):21–??, July 1995. ISSN 0740-1604.

Anonymous:1995:WGPa

[Ano95b] Anonymous. What’s GNU?: Plan 9 (part 1 of 2). Linux Journal, 11:??, March 1995. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic).  URL http://www.
linuxjournal.com/issue11/gnu11.
html.

Anonymous:1995:WGPb

[Ano95c] Anonymous. What’s GNU?: Plan 9 (part 2 of 2). Linux Journal, 12:??, April 1995. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

                                        tronic).  URL http://www.
                                        linuxjournal.com/issue12/gnu12.
                                        html.

Anonymous:1996:CEU

[Ano96a] Anonymous. Cutting edge — use DMI to reduce desktop support costs summer school for IS professionals. Java, Java everywhere. first Plan 9, now Inferno. Go beyond RAID-5. Downsizing of IS slows in 96. How to link relational and object databases. Brace yourself for a price hike for ISDN service. Datamation, 42(11):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:NPS

[Ano96b] Anonymous. A new plan for scalable distributed computing: AT&T’s Plan 9. Data communications, 25(1):106–??, January 1, 1996. CODEN DACODM. ISSN 0363-6399.

Anonymous:1996:POS

[Ano96c] Anonymous. Plan 9: From outer space to kernel space. IEEE Software, 13(3):98–??, May 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1996:PMM

[Ano96d] Anonymous. Plan 9: Multi-media mobile OS. Security, 33(3):64–??, ???? 1996. CODEN SECUEU. ISSN 0890-8826.

Anonymous:19xx:IPD

[Anoxxa] Anonymous. Installing the Plan 9 distribution. Technical report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 19xx. ????

pp. URL http://plan9.att.
com/plan9/doc/install.html.

Anonymous:19xx:P

[Anoxxb] Anonymous. Plan 9. Technical report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 19xx.

???? pp. URL http://plan9.
att.com/plan9/index.html.

Anonymous:19xx:PD

[Anoxxc] Anonymous. Plan 9 distribution. Technical report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 19xx. ????

pp. URL http://plan9.att.
com/plan9/distrib.html.

Anonymous:19xx:PDE

[Anoxxd] Anonymous. Plan 9 distribution errata list. Technical report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 19xx. ????

pp. URL http://plan9.att.
com/plan9/errata.html.

Anonymous:2000:IPD

[Ano00a] Anonymous. Installing the Plan 9 distribution. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000.

URL http://plan9.bell-labs.
com/sys/doc/install.html;
http://plan9.bell-labs.com/
sys/doc/install.pdf;   http:
//plan9.bell-labs.com/sys/
doc/install.ps.

Anonymous:2000:VP

[Ano00b] Anonymous. The various ports. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000.

5 pp. URL http://plan9.
bell-labs.com/sys/doc/port.
html;   http://plan9.bell-
labs.com/sys/doc/port.pdf;
 http://plan9.bell-labs.com/
sys/doc/port.ps. Comments

about the architectures supported by Plan 9.

Bischof:1999:NPK

[B^{+}99] Hans-Peter Bischof et al. Das Netzbetriebssystem Plan 9: Konzepte und Programmierung. Carl Hanser, München, Germany, 1999. ISBN 3-446-18881-9. viii + 242 pp. LCCN ????

Bailey:1996:PUP

[Bai96a] David Bailey. PC UNIX — Plan 9 from AT&T. UNIX review, 14 (1):27–??, ???? 1996. CODEN UNRED5. ISSN 0742-3136.

Bailey:1996:PA

[Bai96b] David Bailey. Plan 9 from AT&T. UNIX review, 14(1):27–28, January 1, 1996. CODEN UNRED5. ISSN 0742-3136.

Ballesteros:2005:TSC

[BGSL05] F. J. Ballesteros, G. Guardiola, E. Soriano, and K. Leal. Traditional systems can work well for pervasive applications. A case study: Plan 9 from Bell Labs becomes ubiquitous. In IEEE [IEE05], pages 295–299. ISBN 0-7695-2300-5 (paperback). LCCN QA76.5915 .I34 2005. URL

http://ieeexplore.ieee.org/
servlet/opac?punumber=9593.

Brazier:1994:DOS

[BJ94] F. M. T. (Frances M. T.) Brazier and D. (Dag) Johansen, editors. Distributed open systems. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-4292-0, 0-8186-4291-2 (microfiche), 0-8186-4290-4 (paperback). LCCN QA76.9.D5 D5563 1994. Includes papers presented at a EurOpen conference held in Tromsø, Norway in 1991, in addition to a number of new papers. IEEE catalog number EH0373-1.

Cox:2002:SP9

[CGP^{+}02] Russ Cox, Eric Grosse, Rob Pike, Dave Presotto, and Sean Quinlan. Security in Plan

9. In USENIX [USE02], pages 3–16. ISBN 1-931971-00-5. LCCN ???? URL http:

//plan9.bell-labs.com/sys/
doc/auth.pdf;   http://www.
usenix.org/publications/library/
proceedings/sec02/cox.html.

Cox:2000:GSP

[Cox00]  Russ Cox.  Getting started

with Plan 9. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. URL http:

                                        //plan9.bell-labs.com/sys/
                                        doc/start.html;   http://plan9.
                                        bell-labs.com/sys/doc/start.
                                        pdf;   http://plan9.bell-labs.
                                        com/sys/doc/start.ps.

Cox:2007:REM

[Cox07] Russ Cox. Regular expression matching can be simple and fast. Report, swtch.com, Cambridge, MA, USA, January 2007.

                                        URL http://swtch.com/~rsc/
                                        regexp/regexp1.html. See also

[Tho68, KP99b, Cox09, Cox10a, Cox12].

Cox:2009:REM

[Cox09] Russ Cox. Regular expression matching: the virtual machine approach. Report, swtch.com, Cambridge, MA, USA, December

                                        2009. URL http://swtch.com/~
                                        rsc/regexp/regexp2.html. See

also [Tho68, KP99b, Cox07, Cox10a, Cox12].

Cox:2010:REM

[Cox10a] Russ Cox. Regular expression matching in the wild. Report, swtch.com, Cambridge, MA, USA, March 2010. URL http://

                                        swtch.com/~rsc/regexp/regexp3.

html. See also [Tho68, KP99b, Cox07, Cox09, Cox12].

Cox:2010:REP

[Cox10b] Russ Cox. re2: an efficient, principled regular expression library. Google Code project, March 2, 2010. URL http://

code.google.com/p/re2/.

Cox:2012:REM

[Cox12] Russ Cox. Regular expression matching with a trigram index, or How Google Code search worked. Report, swtch.com, Cambridge, MA, USA, January

2012. URL http://swtch.com/~
rsc/regexp/regexp4.html. See

also [Tho68, KP99b, Cox07, Cox09, Cox10a].

Deutsch:1967:OE

[DL67] L. Peter Deutsch and Butler W.

Lampson. An online editor. Communications of the ACM, 10(12): 793–799, 803, December 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dorward:1992:P

[DPP92] Sean Dorward, Rob Pike, and Dave Presotto. Plan 9. UNIX review, 10(4):28–??, April 1, 1992. CODEN UNRED5. ISSN 0742-3136.

Duff:1990:RSPa

[Duf90a] T. Duff. Rc — a shell for Plan 9 and Unix systems. EUUG Newsletter, 10(3):12–22, Autumn 1990. CODEN EONLE8. ISSN 1011-4211.

Duff:1990:RSPb

[Duf90b] T. Duff. Rc – a shell for Plan 9 and Unix systems. In UKUUG. UNIX - The Legend Evolves. Proceedings of the Summer 1990 UKUUG Conference, pages 21– 33 (of xi + 260). UK Unix Users Group, Buntingford, Herts, UK, 1990. ISBN 0-9513181-7-9. LCCN ????

Duff:19xx:RSP

[Dufxx] Tom Duff. Rc — a shell for Plan 9 and UNIX systems. Computing Science Technical Report 158g, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 14 pp.

Duff:2000:RPS

[Duf00] Tom Duff. rc — the Plan 9 shell. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 16

pp. URL http://plan9.bell-
labs.com/sys/doc/rc.html;
http://plan9.bell-labs.com/
sys/doc/rc.pdf;   http://plan9.
bell-labs.com/sys/doc/rc.ps.

EurOpen:1991:EUD

[Eur91]  EurOpen, editor.  EurOpen.

UNIX Distributed Open Systems in Perspective. Proceedings of the Spring 1991 EurOpen Conference, Tromsø, Norway, May 20– 24, 1991. EurOpen, Buntingford, Herts, UK, 1991. ISBN 1-873611-00-5. LCCN ????

Fennessy:1996:PBM

[Fen96] Quentin Fennessy. Plan 9: Bad movie, good operating system. Computer, 29(5):117–119, 121– 123, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fillinich:1996:PFF

[Fil96] Paul Fillinich. Plan 9: Feature film to feature-rich OS. BYTE Magazine, 21(3):143–144, March 1, 1996. CODEN BYTEDJ. ISSN 0360-5280.

Flandrena:2000:PM

[Fla00] Bob Flandrena. Plan 9 Mkfiles. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 7 pp.

URL http://plan9.bell-labs.
com/sys/doc/mkfiles.html;
http://plan9.bell-labs.com/
sys/doc/mkfiles.pdf;   http:
//plan9.bell-labs.com/sys/
doc/mkfiles.ps.

Flandrena:2002:AAS

[Fla02] Bob Flandrena. Adding application support for a new architecture in Plan 9. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, May 22, 2002. 7 pp.

URL http://plan9.bell-labs.
com/sys/doc/libmach.pdf.

Ganti:2008:PAL

[Gan08] Ashwin Ganti. Plan 9 authentication in Linux. Operating Systems Review, 42(5):27–33, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Glick:2000:GLP

[Gli00] Paul Glick. Guide to the lp printer spooler. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000.

8 pp. URL http://plan9.bell-
labs.com/sys/doc/lp.html;
http://plan9.bell-labs.com/
sys/doc/lp.pdf;   http://plan9.
bell-labs.com/sys/doc/lp.ps.

Goos:1990:IPS

[Goo90] Anke Goos. Is Plan 9 scifi or UNIX for the future? UNIX/world, 7(10):61–??, October 1, 1990. ISSN 0739-5922.

Gremeyer:1993:UIG

[GS93] Petra Gremeyer and Axel T. Schreiner. It’s ugly but it gets you there: a poor man’s installation guide for Plan 9. Technical report 31, Fachbereich Mathematik/Informatik, Universität Osnabrück, Osnabrück, Germany, 1993. 13 pp.

Hume:2000:MFP

[HF00] Andrew G. Hume and Bob Flandrena. Maintaining files on Plan 9 with Mk. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 14

pp. URL http://plan9.bell-
labs.com/sys/doc/mk.html;

http://plan9.bell-labs.com/
sys/doc/mk.pdf;   http://plan9.
bell-labs.com/sys/doc/mk.ps.

Holzmann:2000:US

[Hol00] Gerard J. Holzmann. Using SPIN. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000.

27 pp. URL http://plan9.
bell-labs.com/sys/doc/spin.
html;   http://plan9.bell-
labs.com/sys/doc/spin.pdf;
 http://plan9.bell-labs.com/
sys/doc/spin.ps.

IEEE:2004:IPD

[IEE04] IEEE, editor. 18th International Parallel and Distributed Processing Symposium: Santa Fe, New Mexico, April 26–30, 2004: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2132-0 (paperback). LCCN QA76.58 .I583 2004. URL

http://ieeexplore.ieee.org/
xpl/RecentCon.jsp?punumber=
9132.

IEEE:2005:TII

[IEE05] IEEE, editor. Third IEEE International Conference on Pervasive Computing and Communications Workshops: proceedings, PerCom 2005 Workshops, 8–12 March, 2005, Kauai Island, Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2300-5 (paperback). LCCN QA76.5915 .I34 2005. URL

                                        http://ieeexplore.ieee.org/
                                        servlet/opac?punumber=9593.

Ierusalimschy:2009:TPM

[Ier09] Roberto Ierusalimschy. A text pattern-matching tool based on Parsing Expression Grammars. Software—Practice and Experience, 39(3):221–258, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jacob:1997:CAM

[JCMS97] B. L. Jacob, P. M. Chen, T. N. Mudge, and S. R. Silverman. A comment on “An analytical model for designing memory hierarchies”. IEEE Transactions on Computers, 46(10):1151, October 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=628401.

See [JCSM96].

Jacob:1996:AMD

[JCSM96] Bruce L. Jacob, Peter M. Chen, Seth R. Silverman, and Trevor N. Mudge. An analytical model for designing memory hierarchies. IEEE Transactions on Computers, 45(10):1180–1194, October 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=543711.

Simulates a storage hierarchy similar to that of Plan 9 [PPTT90b, Qui91]. See comment [JCMS97].

Kernighan:2007:REM

[Ker07] Brian W. Kernighan. A regular expression matcher. In Oram and Wilson [OW07], pages 1– 8. ISBN 0-596-51004-7 (paperback). LCCN QA76.758 .B428 2007; QA76.758 .B43 2007; QA76.758 .B48 2007. URL http:

//www.oreilly.com/catalog/
9780596510046.

Korzeniowski:1996:NPO

[Kor96] Paul Korzeniowski. In the news:

Plan 9: From outer space to kernel space. IEEE Software, 13(3):105, May 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kotsopoulos:1994:PSS

[Kot94] Steve Kotsopoulos. Plan 9 system setup guide (draft). Technical report, University of Toronto, Toronto, ON, Canada, September

9, 1994. URL http://www.ecf.
toronto.edu/plan9/9guide.ps.

Kernighan:1999:PP

[KP99a] Brian W. Kernighan and Rob Pike. The Practice of Programming. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-61586-X. xii + 267 pp. LCCN QA76.6 .K48 1999. US$24.95, CAN$37.50. URL

http://cm.bell-labs.com/cm/
cs/tpop/code.html;   http:
//cseng.aw.com/bookdetail.
qry?ISBN=0-201-61586-X&ptype=
0;   http://tpop.awl.com.

Kernighan:1999:REL

[KP99b] Brian W. Kernighan and Rob Pike. Regular expressions: Languages, algorithms, software. Dr. Dobb’s Journal of Software Tools, 24(4):19–22, April 1999. CO-DEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/ftp/
1999/1999_04/regexp.txt;
http://www.ddj.com/ftp/1999/
1999_04/regexp.zip. See also

[Tho68, Cox07, Cox09, Cox10a, Cox12].

Kuehl:1997:NGP

[Kue97] B. Kuehl. Die naechste Generation — Plan 9. Offene Systeme, 12:16–24, 1997. Proceedings of German UNIX User Group Jahrestagung 13 – September 1997: Wiesbaden, Germany.

Laurikari:2000:NTT

[Lau00] V. Laurikari. NFAs with tagged transitions, their conversion to deterministic automata and application to regular expressions. In Seventh International Symposium on String Processing and Information Retrieval, 2000. SPIRE 2000. 27–29 September 2000, A Curuña, Spain. Proceedings, pages 181–187. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0746-8, 0-7695-0747-6 (case), 0-7695-0748-4 (microfiche). LCCN QA76.9.T48 I59 2000.

Lucent:1997:LPL

[Luc97] Lucent Technologies Inc. The Limbo Programming Language,

1997. URL http://inferno.
lucent.com/inferno/.

Lugosi:1981:POS

[Lug81] Bela Lugosi. Plan 9 from outer space, 1981. 1 videocassette (78 min.).

McIlroy:2004:ESR

[McI04]  M. Douglas McIlroy.  Enu-

merating the strings of regular languages. Journal of Functional Programming, 14(5):503– 518, September 2004. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL http://www.cs.
dartmouth.edu/~doug/nfa.ps.
gz;   https://www.cambridge.
org/core/product/1D46239B6CC6299AA385B3094EBC80E1.

Mirtchovski:2007:WSD

[MI07] Andrey Mirtchovski and Latchesar Ionkov. Why some dead OSes still matter. ;login: the USENIX Association newsletter, 32(5):5–12, October 2007. CO-DEN LOGNEM. ISSN 1044-

6397.  URL https://www.
usenix.org/publications/login/
october-2007-volume-32-number-
5/why-some-dead-oses-still-
matter.

Minnich:2001:PNL

[Min01] Ronald G. Minnich. Private namespaces for Linux. Dr. Dobb’s Journal of Software Tools, 26(12): 23–24, 26, 28, 30, December 2001.

CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Mirtaheri:2014:MPB

[MKM^{+}14] Seyedeh Leili Mirtaheri, Ehsan Mousavi Khaneghah, Amir Saman Memaripour, Lucio Grandinetti, Mohsen Sharifi, and Zarrintaj Bornaee. Multics and Plan 9: The big bangs in the distributed computing system universe. Computing in Science and Engineering, 16(5): 76–85, September/October 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic).  URL http://csdl.
computer.org/csdl/mags/cs/
2014/05/mcs2014050076-abs.
html.

Minnich:2006:XNB

[MM06] R. Minnich and A. Mirtchovski. XCPU: a new, 9p-based, process management system for clusters and grids. In 2006 IEEE International Conference on Cluster Computing: Cluster 2006: Barcelona, September 25th-28th, 2006, pages 1–10. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 1-4244-0328-6. ISSN 1552-5244. LCCN QA76.58.

URL http://ieeexplore.ieee.
org/servlet/opac?punumber=
4100332.

Mirtchovski:2004:PIA

[MS04] A. Mirtchovski and R. Simmonds. Plan 9 — an integrated approach to Grid computing. In IEEE [IEE04], page 273.

ISBN 0-7695-2132-0 (paperback). LCCN QA76.58 .I583 2004. URL

http://ieeexplore.ieee.org/
xpl/RecentCon.jsp?punumber=
9132.

McNaughton:1960:RES

[MY60] R. McNaughton and H. Yamada. Regular expressions and state graphs for automata. IRE Transactions on Electronic Computers, EC-9(1):39–47, March 1960. CODEN IRELAO. ISSN 0367-9950. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5221603.

Ossanna:2000:TUM

[OK00] Joseph F. Ossanna and Brian W.

Kernighan. Troff user’s manual. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 40 pp. URL

http://plan9.bell-labs.com/
sys/doc/troff.html;   http:
//plan9.bell-labs.com/sys/
doc/troff.pdf;   http://plan9.
bell-labs.com/sys/doc/troff.
ps.

Oram:2007:BC

[OW07] Andrew Oram and Greg Wilson, editors. Beautiful code. Theory in practice. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51004-7 (paperback). xxi + 593 pp. LCCN QA76.758 .B428 2007; QA76.758 .B43 2007; QA76.758 .B48 2007.

URL http://www.oreilly.com/
catalog/9780596510046.

Pike:1996:PPMa

[P^{+}96a] Rob Pike et al. Plan 9 Programmer’s Manual: Volume 1: The Manuals. Harcourt Brace and Co., New York, NY, USA, 1996. ISBN ???? ???? pp. LCCN ????

                                        URL http://plan9.bell-labs.
                                        com/plan9/vol1.html.

Pike:1996:PPMb

[P^{+}96b] Rob Pike et al. Plan 9 Programmer’s Manual: Volume 2: The Documents. Harcourt Brace and Co., New York, NY, USA, 1996. ISBN ???? ???? pp. LCCN ????

                                        URL http://plan9.bell-labs.
                                        com/plan9/vol2.html.

Pike:1987:TES

[Pik87] Rob Pike. The text editor sam. Software—Practice and Experience, 17(11):813–845, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pike:1991:PWS

[Pik91] R. Pike. 81/2, the Plan 9 window system. In Proceedings of the Summer 1991 USENIX Conference, Nashville, TN, USA, June 10–14, 1991, pages 257–266 (of x + 473). USENIX, Berkeley, CA, USA, 1991.

Pike:1994:AUI

[Pik94] Rob Pike. Acme: a user interface for programmers. In USENIX Association [USE94], pages 223–234. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994. URL http:

//plan9.bell-labs.com/sys/
doc/acme/acme.html;   http:
//plan9.bell-labs.com/sys/
doc/acme/acme.pdf;   http:
//plan9.bell-labs.com/sys/
doc/acme/acme.ps.

Pike:19xx:PWS

[Pikxxa] Rob Pike. \frac{81}{2}, the Plan 9 window system. Computing Science Technical Report 158d, AT&T Bell Laboratories, Murray Hill, NJ, USA, ?? ??, 19xx. 9 pp.

Pike:19xx:HUP

[Pikxxb] Rob Pike. How to use the Plan 9 C compiler. Web document, 19xx. URL http:

//doc.cat-v.org/plan_9/4th_
edition/papers/comp.

Pike:2000:PWS

[Pik00a] Rob Pike. 81/2, the Plan 9 window system. Technical report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 10 pp. URL http:

//plan9.bell-labs.com/sys/
doc/8%bd/8%bd.html;   http:
//plan9.bell-labs.com/sys/
doc/8%bd/8%bd.pdf;   http:
//plan9.bell-labs.com/sys/
doc/8%bd/8%bd.ps. World-Wide

Web document. Originally appeared in [Pik91].

Pike:2000:HUP

[Pik00b] Rob Pike. How to use the Plan 9 C compiler. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000.

14 pp. URL http://plan9.
bell-labs.com/sys/doc/comp.
html;   http://plan9.bell-
labs.com/sys/doc/comp.pdf;
 http://plan9.bell-labs.com/
sys/doc/comp.ps.

Pike:2000:LFNa

[Pik00c] Rob Pike. Lexical file names in Plan 9, or getting dot-dot right. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 11 pp. URL http:

//plan9.bell-labs.com/sys/
doc/lexnames.html;   http:
//plan9.bell-labs.com/sys/
doc/lexnames.pdf;   http://
plan9.bell-labs.com/sys/doc/
lexnames.ps;   https://9p.
io/sys/doc/lexnames.html.

Pike:2000:LFNb

[Pik00d] Rob Pike. Lexical file names in Plan 9, or getting dot-dot right. In USENIX [USE00], pages 85– 92. ISBN 1-880446-22-7.

Pike:2000:MPA

[Pik00e] Rob Pike. A manual for the plan 9 assembler. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000.

10 pp. URL http://plan9.
bell-labs.com/sys/doc/asm.
html;   http://plan9.bell-
labs.com/sys/doc/asm.pdf;
http://plan9.bell-labs.com/
sys/doc/asm.ps.

Pike:2000:POU

[Pik00f]  Rob Pike.  Plumbing and

other utilities. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 16 pp. URL

http://plan9.bell-labs.com/
sys/doc/plumb.html;   http:
//plan9.bell-labs.com/sys/
doc/plumb.pdf;   http://plan9.
bell-labs.com/sys/doc/plumb.
ps.

Pike:2000:SSR

[Pik00g] Rob Pike. System software research is irrelevant. Web talk slides., February 21, 2000.

URL http://herpolhode.com/
rob/utah2000.pdf.

Pike:2000:TES

[Pik00h] Rob Pike. The text editor sam. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, June 7, 2000. 30 pp. URL

http://plan9.bell-labs.com/
sys/doc/sam/sam.pdf.

Pike:2002:CPE

[Pik02] Rob Pike. Changes to the programming environment in the fourth release of Plan 9. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, May 22, 2002. 6 pp.

URL http://plan9.bell-labs.
com/sys/doc/prog4.pdf.

Pike:1995:PBL

[PPD^{+}95] Rob Pike, Dave Presotto, Sean Dorward, Bob Flandrena, Ken Thompson, Howard Trickey, and Phil Winterbottom. Plan 9 from Bell Labs. Computing Systems, 8 (3):221–254, Summer 1995. CO-DEN CMSYE2. ISSN 0895-6340.

URL http://plan9.bell-labs.
com/sys/doc/9.html;   http:
//plan9.bell-labs.com/sys/
doc/9.pdf;   http://plan9.
bell-labs.com/sys/doc/9.ps.

Pike:1991:DPC

[PPT91a] Rob Pike, Dave Presotto, and Ken Thompson. Designing Plan 9: contrary to what you might expect, tomorrow’s distributed computer systems may look a lot like Plan 9. Dr. Dobb’s Journal of Software Tools, 16(1):49–??, January 1, 1991. CODEN DDJOEB. ISSN 1044-789X.

Pike:1991:PEP

[PPT^{+}91b] Rob Pike, Dave Presotto, Ken Thompson, Howard Trickey, Tom Duff, and Gerard Holzmann. Plan 9: The early papers. Computing Science Technical Report 158, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1991. i + 55 pp.

URL https://telecomarchive.
s3.us-east-2.amazonaws.com/
docs/bsp-archive/Letters%20and%
20Memos/CSTR/CSTR%20158.pdf.

Pike:1992:UNS

[PPT^{+}92] Rob Pike, Dave Presotto, Ken Thompson, Howard Trickey, and Phil Winterbottom. The use of name spaces in Plan 9. In Proceedings of the 5th workshop on ACM SIGOPS European workshop: Models and paradigms for distributed systems structuring, pages 1–5. ACM Press, New York, NY 10036, USA, 1992.

URL http://www.informatik.
hu-berlin.de/~mint/Library/
Plan9/names.html.

Pike:1993:UNS

[PPT^{+}93] Rob Pike, Dave Presotto, Ken Thompson, Howard Trickey, and Phil Winterbottom. The use of name spaces in Plan 9. Operating Systems Review, 27(2):72– 76, April 1993. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL

http://plan9.bell-labs.com/
sys/doc/names.html;   http:
//plan9.bell-labs.com/sys/
doc/names.pdf;   http://plan9.
bell-labs.com/sys/doc/names.
ps.

Pike:19xx:PEP

[PPT^{+}xx] Rob Pike, Dave Presotto, Ken Thompson, Howard Trickey, Tom Duff, and Gerard Holzmann. Plan 9: The early papers. Computing Science Technical Report 158a, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 1 pp.

Pike:19xx:PSW

[PPTHxx] Rob Pike, Dave Presotto, Ken Thompson, and Gerard Holzmann. Process sleep and wakeup on a shared-memory multiprocessor. Computing Science Technical Report 158f, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 6 pp.

Pike:2000:PSW

[PPTH00] Rob Pike, Dave Presotto, Ken Thompson, and Gerard Holzmann. Process sleep and wakeup on a shared-memory multiprocessor. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ,

USA, 2000.  6 pp.  URL

                                        http://plan9.bell-labs.com/
                                        sys/doc/sleep.html;   http:
                                        //plan9.bell-labs.com/sys/
                                        doc/sleep.pdf. Appeared in

a slightly different form in Proceedings of the Spring 1991 EurOpen Conference, Tromsø, Norway, 1991, pp. 161–166.

Pike:1990:PBLa

[PPTT90a] R. Pike, D. Presotto, K. Thompson, and H. Trickey. Plan 9 from Bell Labs. EUUG Newsletter, 10 (3):2–11, Autumn 1990. CODEN EONLE8. ISSN 1011-4211.

Pike:1990:PBLb

[PPTT90b] R. Pike, D. Presotto, K. Thompson, and H. Trickey. Plan 9 from Bell Labs. In UKUUG. UNIX - The Legend Evolves. Proceedings of the Summer 1990 UKUUG Conference, pages 1–9 (of xi + 260). UK Unix Users Group, Buntingford, Herts, UK, 1990. ISBN 0-9513181-7-9. LCCN ???? See also [JCSM96].

Pike:1991:DP

[PPTT91a] Rob Pike, Dave Presotto, Ken Thompson, and Howard Trickey. Designing Plan 9. Dr. Dobb’s Journal of Software Tools, 16(1): 49–50, 52, 54, 56–60, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Presotto:1991:PDS

[PPTT91b] D. Presotto, R. Pike, K. Thompson, and H. Trickey. Plan 9, a distributed system. In EurOpen [Eur91], pages 43–50. ISBN 1-873611-00-5. LCCN ????

Presotto:1992:PDS

[PPTT92] D. Presotto, R. Pike, K. Thompson, and H. Trickey. Plan 9, a distributed system. In Proceedings of the USENIX Workshop on Micro-Kernels and Other Kernel Architectures, pages 31–37 (of 303). USENIX, Berkeley, CA, USA, 1992.

Presotto:1994:PDS

[PPTT94] D. Presotto, R. Pike, K. Thompson, and H. Trickey. Plan 9: a distributed system. In Brazier and Johansen [BJ94], pages 49– 56. ISBN 0-8186-4292-0, 0-8186-4291-2 (microfiche), 0-8186-4290-4 (paperback). LCCN QA76.9.D5 D5563 1994. Includes papers presented at a EurOpen conference held in Tromsø, Norway in 1991, in addition to a number of new papers. IEEE catalog number EH0373-1.

Pike:19xx:PBL

[PPTTxxa] Rob Pike, Dave Presotto, Ken Thompson, and Howard Trickey. Plan 9 from Bell Labs. Computing Science Technical Report 158b, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 9 pp.

Presotto:19xx:PDS

[PPTTxxb] Dave Presotto, Rob Pike, Ken Thompson, and Howard Trickey. Plan 9, a distributed system. Computing Science Technical Report 158c, AT&T Bell Laboratories, Murray Hill, NJ, USA, ?? ??, 19xx. 7 pp.

Presotto:1988:PBL

[Pre88] David Leo Presotto. Plan 9 from Bell Labs — the network. In USENIX Association [USE88], pages 15–21.

Presotto:1990:MSP

[Pre90] D. L. Presotto. Multiprocessor streams for Plan 9. In UKUUG. UNIX - The Legend Evolves. Proceedings of the Summer 1990 UKUUG Conference, pages 11– 19 (of xi + 260). UK Unix Users Group, Buntingford, Herts, UK, ???? 1990. ISBN 0-9513181-7-9. LCCN ????

Presotto:1993:MSPa

[Pre93a] D. Presotto. Multiprocessor STREAMS for Plan 9. EUUG Newsletter, ??(??):??, January 1993. CODEN EONLE8. ISSN 1011-4211.

Presotto:1993:MSPb

[Pre93b] David Presotto. Multiprocessor streams for Plan 9. In USENIX Association. Proceedings of the Winter 1993 USENIX Conference, page ?? USENIX, Berkeley, CA, USA, 1993.

Presotto:19xx:MSP

[Prexx] David Leo Presotto. Multiprocessor streams for Plan 9. Computing Science Technical Report 158e, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 10 pp.

Pike:1993:HW

[PT93] Rob Pike and Ken Thompson.

Hello world. In USENIX Association [USE93], pages 43–50. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993. URL

http://plan9.bell-labs.com/
sys/doc/utf.pdf.

Pike:2000:HWL

[PT00] Rob Pike and Ken Thompson.

Hello world or Kαλημ ´ ϵρα κ oσμϵ or …. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA,

2000.  URL http://plan9.
bell-labs.com/sys/doc/utf.
html;   http://plan9.bell-
labs.com/sys/doc/utf.pdf;
http://plan9.bell-labs.com/
sys/doc/utf.ps. Originally ap-

peared, in a slightly different form, in Proc. of the Winter 1993 USENIX Conf., pp. 43–50, San Diego.

Presotto:1993:ONP

[PW93] D. Presotto and P. Winterbottom. The organization of networks in Plan 9. In USENIX Association. Proceedings of the Winter 1993 USENIX Conference, pages 271– 280 (of x + 530). USENIX, Berkeley, CA, USA, 1993.

URL http://plan9.bell-labs.
com/sys/doc/net/net.html;
http://plan9.bell-labs.com/
sys/doc/net/net.pdf;   http:
//plan9.bell-labs.com/sys/
doc/net/net.ps.

Presotto:2000:P

[PW00] Dave Presotto and Phil Winterbottom. The IL protocol. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 5 pp. URL

http://doc.cat-v.org/plan_
9/4th_edition/papers/il/;
http://plan9.bell-labs.com/
sys/doc/il/il.html;   http:
//plan9.bell-labs.com/sys/
doc/il/il.pdf;   http://plan9.
bell-labs.com/sys/doc/il/il.
ps.

Quinlan:2002:VNA

[QD02] Sean Quinlan and Sean Dorward. Venti: a new approach to archival storage. Technical report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, May 2, 2002. 13 pp.

URL http://plan9.bell-labs.
com/sys/doc/venti/venti.pdf.

Quinlan:1991:CWF

[Qui91] S. Quinlan. A cached WORM file system. Software—Practice and Experience, 21(12):1289–??, December 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See also [JCSM96].

Rayome:1994:SSP

[Ray94] Jerry Kent Rayome. System security of the Plan 9 distributed operating system from Bell Laboratories. Thesis (m.s.), California State University, Chico, Chico, CA, USA, 1994. viii + 74 pp.

Ritchie:19xx:IHQ

[Ritxx] Dennis Ritchie. An incomplete history of the QED text editor. Report, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 19xx.

URL http://plan9.bell-labs.
com/who/dmr/qed.html.

Robbins:1995:WGPa

[Rob95a] Arnold Robbins. What’s GNU?

[Plan 9 Part I]. Linux Journal, 10:??, March 1995. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

http://interactive.linuxjournal.
com/article/1012.

Robbins:1995:WGPb

[Rob95b] Arnold Robbins. What’s GNU?

Plan 9 Part II. Linux Journal, 10: ??, April 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//interactive.linuxjournal.

                                        com/article/0062;   http://
                                        portal.acm.org/citation.cfm?
                                        id=324810.

Rabin:1959:FAT

[RS59] M. O. Rabin and D. Scott. Finite automata and their decision problems. IBM Journal of Research and Development, 3(??): 114–125, 1959. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Skocovski:1998:UPP

[Sko98] Ludek Skocovski. UNIX, POSIX, Plan 9. ????, Brno, Czech Republic, 1998. ISBN 80-902612-0-5. 394 pp. In Czech.

Stallman:2012:PEP

[Sta12] Richard Stallman. The problems of the (earlier) Plan 9 license. Web essay., June 10,

                                        2012. URL http://www.gnu.
                                        org/philosophy/plan-nine.html.

See [MI07] for comments on this essay, and how its objections were resolved in newer releases of Plan 9.

Thompson:19xx:BSP

[TCxx] Ken Thompson and Geoff Collyer. The 64-bit standalone Plan 9 file server. Web site., 19xx. URL

                                        http://doc.cat-v.org/plan_
                                        9/4th_edition/papers/fs/.

Thompson:1968:PTR

[Tho68] Ken Thompson. Programming techniques: Regular expression search algorithm. Communications of the ACM, 11(6):419–422, June 1968. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//patft.uspto.gov/. See also

[KP99b, Cox07, Cox09, Cox10a, Cox12].

Thompson:19xx:NCC

[Thoxxa] Ken Thompson. A new C compiler. Computing Science Technical Report 158h, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 12 pp. URL

http://doc.cat-v.org/bell_
labs/new_c_compilers/new_c_
compiler.pdf.

Thompson:19xx:PCC

[Thoxxb] Ken Thompson. Plan 9 C compilers. Web site, 19xx. URL http:

//doc.cat-v.org/plan_9/4th_
edition/papers/compiler.

Thompson:2000:PCC

[Tho00a] Ken Thompson. Plan 9 C compilers. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 11 pp. URL http:

//plan9.bell-labs.com/sys/
doc/compiler.html;   http:
//plan9.bell-labs.com/sys/
doc/compiler.pdf;   http://
plan9.bell-labs.com/sys/doc/

compiler.ps. Originally appeared, in a different form, in Proceedings of the Summer 1990 UKUUG Conference, pp. 41–51, London, 1990.

Thompson:2000:PFS

[Tho00b] Ken Thompson. The Plan 9 file server. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ,

USA, 2000.  7 pp.  URL

                                        http://plan9.bell-labs.com/
                                        sys/doc/fs/fs.html;   http:
                                        //plan9.bell-labs.com/sys/
                                        doc/fs/fs.pdf;   http://plan9.
                                        bell-labs.com/sys/doc/fs/fs.
                                        ps.

Trbovich:2016:ICB

[Trb16] Pete Trbovich. It came from Bell Labs. Web site., November 15, 2016.

Trickey:2000:AAP

[Tri00]  Howard Trickey.  APE —

the ANSI/POSIX environment. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 4

                                        pp.  URL http://plan9.
                                        bell-labs.com/sys/doc/ape.
                                        html;   http://plan9.bell-
                                        labs.com/sys/doc/ape.pdf;
                                        http://plan9.bell-labs.com/
                                        sys/doc/ape.ps.

USENIX:1988:ECP

[USE88] USENIX Association, editor. EUUG Conference Proceedings, Spring, 1988. London, England. EUUG, Buntingford, Herts, UK, Spring 1988.

USENIX:1992:PUW

[USE92] USENIX, editor. Proceedings of the USENIX Workshop on Micro-Kernels and Other Kernel Architectures: 27–28 April, 1992, Seattle, WA, USA. USENIX, Berkeley, CA, USA, 1992. ISBN 1-880446-42-1. LCCN QAX 32.

USENIX:1993:PWU

[USE93] USENIX Association, editor.

Proceedings of the Winter 1993 USENIX Conference: January 25–29, 1993, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

USENIX:1994:PWU

[USE94] USENIX Association, editor.

Proceedings of the Winter 1994 USENIX Conference: January 17–21, 1994, San Francisco, California, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

USENIX:2000:UAT

[USE00] USENIX, editor. 2000 USENIX Annual Technical Conference: San Diego, CA, USA, June 18– 23, 2000. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-22-7.

USENIX:2002:PUS

[USE02] USENIX, editor. Proceedings of the 11th USENIX Security Symposium 2002, August 5–9, 2002, San Francisco CA. USENIX, Berkeley, CA, USA, 2002. ISBN 1-931971-00-5. LCCN ????

VanHensbergen:2008:HAR

[VFMM08] Eric Van Hensbergen, Charles Forsyth, Jim McKie, and Ron Minnich. Holistic aggregate resource environment. Operating Systems Review, 42(1):85– 91, January 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Welch:1993:CTD

[Wel93] Brent Welch. A comparison of three distributed file system architectures: Vnode, Sprite, and Plan 9. Technical report CSL-93-18, Xerox Corp., Palo Alto Research Center, Palo Alto, CA, USA, December 1993. 18 pp.

Welch:1994:CTD

[Wel94] B. Welch. A comparison of three distributed file system architectures: Vnode, Sprite, and Plan 9. Computing Systems, 7(2):175– 199, Spring 1994. CODEN CM-SYE2. ISSN 0895-6340.

Winterbottom:1994:ADB

[Win94] Philip Winterbottom. ACID: a debugger based on a language. In USENIX Association [USE94], page ?? ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994. URL http:

//plan9.bell-labs.com/sys/
doc/acidpaper.html;   http:
//plan9.bell-labs.com/sys/
doc/acidpaper.pdf;   http:
//plan9.bell-labs.com/sys/
doc/acidpaper.ps.

Winterbottom:2000:ADB

[Win00a] Phil Winterbottom. Acid: a debugger built from a language. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000. 14 pp. URL http:

//plan9.bell-labs.com/sys/
doc/acidpaper.html;   http:
//plan9.bell-labs.com/sys/
doc/acidpaper.pdf;   http:
//plan9.bell-labs.com/sys/
doc/acidpaper.ps. Originally

appeared in Proc. of the Winter 1994 USENIX Conf., pp. 211– 222, San Francisco, CA.

Winterbottom:2000:AM

[Win00b] Phil Winterbottom. Acid manual. World-Wide Web document, Computing Sciences Research Center, Bell Laboratories, Murray Hill, NJ, USA, 2000.

25 pp. URL http://plan9.
bell-labs.com/sys/doc/acid.
html;   http://plan9.bell-
labs.com/sys/doc/acid.pdf;
 http://plan9.bell-labs.com/
sys/doc/acid.ps.

Wood:1984:POS

[WJV^{+}84] Edward D. (Edward Davis)

Wood, Tor Johnson, Vampira, Tom Keene, and Gregory Walcott. Plan 9 from outer space, 1984. ISBN 1-55511-689-2. 1 videocassette (approx. 79 min.).

Zobel:1993:SLL

[ZMSD93] Justin Zobel, Alistair Moffat, and Ron Sacks-Davis. Searching large lexicons for partially specified terms using compressed inverted files. In Agrawal et al. [ABB93], pages 290–301. ISBN 1-55860-152-X. LCCN QA76.9.D3

I61 1993. URL http://www.
vldb.org/dblp/db/conf/vldb/
ZobelMS93.html. Co-sponsored

by VLDB Endowment and Irish Computer Society; in cooperation with the IEEE Technical Committee on Data Engineering.